### v1

# Re-Examine Distantly Supervised NER: A New Benchmark and a Simple Approach

Abstract
This paper delves into Named Entity Recognition (NER) under the framework of Distant Supervision (DS-NER), where the main challenge lies in the compromised quality of labels due to inherent errors such as false positives, false negatives, and positive type errors. We critically assess the efficacy of current DS-NER methodologies using a real-world benchmark dataset named QTL, revealing that their performance often does not meet expectations. To tackle the prevalent issue of label noise, we introduce a simple yet effective approach, Curriculum-based Positive-Unlabeled Learning (CuPUL), which strategically starts on “easy” and cleaner samples during the training process to enhance model resilience to noisy samples. Our empirical results highlight the capability of CuPUL to significantly reduce the impact of noisy labels and outperform existing methods.
Re-Examine Distantly Supervised NER:
A New Benchmark and a Simple Approach
Yuepei Li                        Kang Zhou                        Qiao Qiao Iowa State University {liyp0095, kangzhou, qqiao1, qingwang, qli}@iastate.edu                        Qing Wang                        Qi Li
1 Introduction
Distantly-Supervised Named Entity Recognition (DS-NER) is a task to leverage existing knowledge bases (KBs) or dictionaries to provide annotations for named entity recognition tasks. This approach significantly reduces the need for labor-intensive manual annotations, but it faces challenges due to issues in automated annotations, such as false positives, false negatives, and positive type errors. To address the annotation errors, various methods are proposed. Some studies focus on false negative issues Shang et al. (2018); Peng et al. (2019); Zhou et al. (2022). Others propose to tackle general noisy annotations through noise removal processes Meng et al. (2021); Liang et al. (2020); Hedderich and Klakow (2018); Zhang et al. (2021a); Liu et al. (2021).
DS-NER approaches have been successfully applied to existing NER benchmark datasets, achieving competitive performances with fully supervised methods. However, in our usage in a real-life DS-NER application, we observe that the existing DS-NER approaches fail to obtain satisfactory performances. In this paper, we present a new benchmark dataset named QTL and re-examine the existing DS-NER methods in a real-life DS-NER task. We further present a simple yet effective approach following the philosophy of curriculum learning.
Previous DS-NER approaches primarily rely on noise removal to address annotation issues especially false positives and positive type errors. However, these methods face several challenges when applied in real-life DS-NER tasks. First, a moderately well-trained NER model is necessary to detect noise. However, it is hard to determine proper hyper-parameters when an NER model is moderately well-trained without sufficient validation data. Stopping the training too early, the model cannot produce an accurate enough model for noise detection. Stopping the training too late, the model will learn the noise and degrade the performance. Second, the moderately well-trained NER model is trained on noisy labels initially, so the noise detection methods may have unknown biases and cause irreparable damage.
Instead of removing noisy labels, we propose a simple yet robust approach using token-level Curriculum-based Positive-Unlabeled Learning (CuPUL) to tackle the challenge of noisy labels in DS-NER tasks. The motivation of curriculum learning is that deep learning models are non-convex and trained using batches of samples, so the order of training data can significantly impact the model performance. Curriculum Learning rearranges the batches of training samples such that the model learns from easy to hard and learns from easy samples more times. With the new arrangement, the models tend to converge to a better local optimum. We further design a token-level curricula arrangement to address the token-level noise for DS-NER tasks. We observe that “easy samples” are usually cleaner. Consequentially, learning from easy samples first can avoid label noise initially and make the model more robust. We adopt the Positive-Unlabeled (PU) learning paradigm to address the false negative issues.
In summary, our main contributions are:
  • We present a real-life DS-NER application, QTL, and test the performance of the existing state-of-the-art methods. We observe that many methods do not follow the practical DS-NER setting and have unsatisfactory performance.
  • We propose a simple method CuPUL to address the noise issue in DS-NER. We empirically demonstrate that CuPUL can significantly outperform the state-of-the-art DS-NER method on the QTL dataset.
  • CuPUL also achieves good performance on existing benchmarks.
2 Related Work
2.1 Benchmark Datasets
To reduce the cost of human-annotated training data for NER tasks, DS-NER uses professional dictionaries or knowledgebases for annotations. Existing DS-NER benchmark datasets use NER benchmark datasets to simulate the distant supervision setting by replacing the human annotations on training datasets with knowledgebase annotations Liang et al. (2020); Shang et al. (2018); Zhou et al. (2022). There are some potential biases of current DS-NER benchmarks. 1) only BC5CDR Shang et al. (2018) dataset comes from professional domains where DS-NER tasks are in high demand. 2) a series of hand-crafted procedures were applied to the current DS-NER benchmarks. Such procedures are entity-type dependent and require substantial human effort, and thus may not be generalizable to other DS-NER tasks. 3) The major entity types in existing DS-NER benchmarks consist of many proper nouns (such as Person’s names, location, and gene names), but in many applications, the entities to recognize need not to be proper nouns.
2.2 DS-NER Methods
Handling annotation errors in DS-NER tasks has drawn special attention. Here we briefly discuss a few representative approaches.
One line of work assumes that distant supervision often has high-quality positive labels, therefore focusing on alleviating the impact of false negative errors. Some methods address this issue by detecting potential entity candidates Shang et al. (2018); Xu et al. (2023). Some methods adopt positive and unlabeled learning to tackle false negative errors from the loss estimation perspective without separate detection steps Peng et al. (2019); Zhou et al. (2022). Due to its superiority in tolerating false negative errors, we embed Conf-MPU Zhou et al. (2022) into our proposed method.
Another line of work simultaneously considers annotation errors of all types. Some methods propose to train an initial model and apply a self-training framework to reduce the impact of noise Liang et al. (2020); Liu et al. (2021); Zhang et al. (2021b); Qu et al. (2023); Li et al. (2023). Methods such as RoSTER Meng et al. (2021) and SANTA Si et al. (2023) endeavor to diminish noise effects using loss functions tailored for noise resilience.
These DS-NER approaches obtain promising performance on existing DS-NER benchmark datasets. However, in a real-life DS-NER application, we observe that they fail to obtain satisfactory performances. These methods are trained on noisy labels initially, so the noise detection may have unknown biases and cause irreparable damage.
3 QTL Benchmark
We present QTL, a real-life DS-NER application in the animal science domain. The entity type to recognize is “trait”, an important task in the construction of genotype-phenotype databases for advancing livestock genomics research and breeding methodologies. Different from previous DS-NER datasets, where entities consist of many proper nouns, trait entities consist of descriptive expressions. For example, in the sentence “There are several reports on candidate gene analysis of subcutaneous fat depth traits using similar Wagyu × Limousin F2 populations”, “subcutaneous fat depth traits” is a trait entity that needs to be identified.
We collected a corpus with 1,716 abstracts from quantitative trait locus (QTL) studies related to six species, which were meticulously selected from PubMed [^1] https://pubmed.ncbi.nlm.nih.gov/ by domain experts. There are 18706 sentences and 514176 tokens in the train set. For the distant annotation process, we gathered a specialized dictionary of 3,884 trait names from four established domain ontologies [^2] Vertebrate Trait (VT) Ontology, Livestock Product Trait (LPT) Ontology, Livestock Breed Ontology (LBO), and Clinical Measurement Ontology (CMO). To evaluate DS-NER approaches, a well-trained domain curator provided ground truth annotations for 107 abstracts. These were then divided into validation and test sets, containing 21 and 1,044 sentences, 952 tokens and 32251 tokens, respectively. The test set has 1219 entities.
It is noteworthy that in the real-life application of QTL, the validation set used for hyperparameter tuning is quite small, comprising only 21 sentences. This limited size of the validation set may impact the tuning of hyperparameters during the model training process, potentially affecting the model’s performance. This issue reflects a realistic challenge encountered in DS-NER applications in real-life scenarios.
Metrics: In the QTL application, according to the curator’s practical needs, identifying potential entities is more important than identifying precise boundaries. Therefore, besides the strict span-level Precision (P), Recall (R), and F1 scores, we also introduce relaxed Precision (P), Recall (R), and F1 scores to evaluate the performance of DS-NER methods for practical usage. For relaxed metrics, it deems a predicted span correct if there is at least one overlapping word with the ground truth annotation.
4 Methodology
Figure 1: The CuPUL Framework
In DS-NER, we denote an input sentence with $M$ tokens as $\boldsymbol{x}=[x_{1},x_{2},\cdots,x_{M}]$ and denote corresponding annotations as $\boldsymbol{y}=[y_{1},y_{2},\cdots,y_{M}]$, $y_{i}\in\{0,1,\cdots,k\}$, where 0 denotes unlabeled types and $1,\cdots,k$ denote $k$ entity types. For the models, a pre-trained language model such as RoBERTa is used to encode token representations and it is followed by a softmax function to forward the prediction of entity labels for each token in the sentence.
Curriculum learning has two main steps: difficulty estimation and curriculum scheduler Kocmi and Bojar (2017). We have more discussion and related work in Appendix .1. Following the main steps of curriculum learning, we design the framework of CuPUL, which is shown in 1. The process starts by distantly labeling the corpus using knowledge bases, and then several voters are trained using this data to calculate token difficulty scores. Finally, CuPUL trains a NER classifier following the curriculum scheduler using confidence-based positive-unlabeled learning.
4.1 Difficulty Estimation
Motivated by the token-level noises in DS-NER tasks, we design the difficulty estimator and the curriculum scheduler at the token level as well. It allows the model to learn from one sentence by ignoring the noisy tokens. For example, in the sentence “Peter(PER) lives(O) in(O) America(ORG)”, “Peter”, “lives”, and “in” are clean samples, and “America” is a noise sample. The model can learn from “Peter lives in X” by ignoring the noise in the sentence. The token’s difficulty score reflects its inherent learnability. These scores are estimated using the disagreements between basic NER models or voters.
4.1.1 Voters
The design of the voters demands simplicity and variability. To balance efficiency and diversity, several voters are trained to increase prediction results variability. Label imbalance in NER tasks is mitigated by sampling from negative samples, introducing randomness to voters. Randomness is further enhanced with different random processes, such as shuffling and initialization.
For training the voters, a neural network for NER classification is used, as defined at the beginning of Section 4. The voters are trained using a regular multi-class classification risk function. The training process follows the Positive-Negative setting, where 0 represents non-entity type. This is doable because the performance of the voter itself does not affect the final outcome of CuPUL, which we will introduce in the section 4.2.
4.1.2 Difficulty Scores
After training $V$ voters, each token $x$ receives $V$ predicted class probabilities $f(x,\boldsymbol{\theta}_{1}),...,f(x,\boldsymbol{\theta}_{V})$, where $\boldsymbol{\theta}_{1}...\boldsymbol{\theta}_{V}$ are the voters’ parameters. The prediction $f(x,\boldsymbol{\theta}_{i})$ is a vector that represents the class distribution of each token $x$ denoted as $\boldsymbol{\Pr}_{i}(x)$. The difficulty of the token is assessed based on the disagreement among these distributions. Specifically, we use Kullback-Leibler (KL) divergence, a measurement for dissimilarities of two distributions $\boldsymbol{\Pr}_{i}(x)$ and $\boldsymbol{\Pr}_{j}(x)$, to calculate the disagreement level of two voters. Mathematically, it is:
$H_{ij}=\frac{1}{2}\{D_{KL}(\boldsymbol{\Pr}_{i}(x)||\boldsymbol{\Pr}_{j}(x)))+\\ D_{KL}(\boldsymbol{\Pr}_{j}(x))||\boldsymbol{\Pr}_{i}(x))\},$ (1)
where $D_{KL}(\cdot)$ denotes the KL divergence. KL divergence is asymmetric. By taking the average of $H_{ij}$ and $H_{ji}$, we derive a symmetric difficulty score $H_{\{ij\}}$.
Given that there are $V$ voters, the final difficulty score for each token $x$ is defined as the average of the non-identical pairs among all voters as:
$H=\frac{\sum_{i=1}^{V}\sum_{j=i+1}^{V}H_{\{ij\}}}{V\cdot(V-1)/2}.$ (2)
Eq.(2) defines the token difficulty scores as an arithmetic mean of disagreements between pair-wise voters. Consequently, a token’s difficulty score is low when all voters agree, and it increases with greater disagreement.
4.2 Curriculum Design
To avoid overfitting of negative samples, we set the model learning process as Positive-Unlabeled (PU) learning, where data labeled with 0 is considered unlabeled rather than non-entity. PU learning assumes that the unlabeled data follows the distribution of the entire dataset Zhou et al. (2022). To fulfill the assumption, we directly incorporate all unlabeled data into the curriculum to prevent too less unlabeled tokens in curricula. Different curriculum partitions are executed solely on the labeled positive data. In this way, only the learning order of positive samples can affect CuPUL performance.
Our curriculum is built on token difficulty scores $H$, which are mostly "easy" as they follow a long-tail distribution (Figure 3). Previous work Platanios et al. (2019); Gnana Sheela and Deepa (2013) suggests that a curriculum with uniformly ranged difficulty scores might lead to most tokens belonging to the first curriculum, making curriculum learning ineffective. Hence, we propose using a power-law selector to construct a more effective curriculum scheduler.
To build the curricula, we first arrange $T_{u}$ unlabeled tokens followed by $T_{p}$ positive-labeled tokens sorted by their difficulty scores in ascending order. The first curriculum consists of all unlabeled tokens and the first $\tau T_{p}$ labeled positive tokens, where $\tau$ ($10<\tau<$ 1) is a selective factor. The second curriculum consists of the first $\tau^{2}T$ tokens from the remaining $(1-\tau)T_{p}$ tokens. This selection process continues until the penultimate curriculum. The remaining tokens are placed in the final curriculum. These curricula are denoted as $C_{1},C_{2},...,C_{\eta}$:
$\begin{split}C_{1}&:1\sim T_{u}+\tau T_{p}\\ C_{2}&:T_{u}+\tau T_{p}+1\sim T_{u}+(\tau+\tau^{2})T_{p}\\ &\cdots\\ C_{\eta}&:T_{u}+(\tau^{\eta-1}+\cdots+\tau)T_{p}+1\sim T_{u}+T_{p}.\\ \end{split}$
For example, suppose $T_{p}=20$, $T_{u}=80$, $\tau=0.5$, and $\eta=3$. Then, the first curriculum consists of tokens indexed from 1 to 90 ($80$ unlabeled tokens and the $10$ easiest positive tokens), the second curriculum consists of tokens indexed from 91 to 95, and the last curriculum consists of tokens indexed from 96 to 100.
4.3 Curriculum-based PU Learning
We train the NER classifier across $\eta$ curricula using the “Baby Step” training scheduleSpitkovsky et al. (2010); Cirik et al. (2017). Starting with $C_{1}$, we add each subsequent curriculum after a fixed number of epochs, training through all curricula until completion. The training stages ($\{S_{i},1<i\leq\eta\}$) correspond to the number of curricula, with the model trained over multiple epochs in each stage. Each stage is treated as an independent training segment, with earlier curricula being reviewed more frequently, enhancing learning under PU assumptions and resulting in a robust curriculum learning framework.
Specifically, we adopt the Conf-MPU loss function, proposed by Zhou et al. (2022), as the backbone PU loss function in the curriculum-based training. Conf-MPU loss function has been shown to be more robust to PU assumption violation in practice. Instead of having entity confidence score $\lambda(x)$ estimated by another binary PU model, the only difference we make is to reuse the voters trained in Section 4.1 to ensemble the confidence score for each token $x$. We use the soft-label ensemble as
$\boldsymbol{\Pr}(x)=\frac{\sum_{j=1}^{V}f(x,\boldsymbol{\theta}_{j})}{V},$ (3)
where $\boldsymbol{\Pr}(x)$ is the ensemble probability distribution over all classes.
The confidence score of a token $x$ being an entity token is then calculated as
$\lambda(x)=\sum_{j=1}^{k}\boldsymbol{\Pr}_{j}(x).$ (4)
For the neural network of the NER classifier, we choose the same structure with voters, which is defined at the beginning of Section 4.
4.4 Self-Training
Several studies Liang et al. (2020); Peng et al. (2019); Meng et al. (2021) have shown that self-training can effectively upgrade the performance of a trained DS-NER model. We apply the self-training method in Meng et al. (2021), which uses soft labels to conduct self-training and uses a masked language model to conduct contextual data augmentation simultaneously. Self-training is used directly after CuPUL, and we call the final classifier after self-training CuPUL+ST.
5 Experimental Study
5.1 Baseline Methods
We use two groups of baseline methods.
Fully supervised methods. We include a fully supervised NER method based on the RoBERTa-base model Liu et al. (2019) as an upper-bound performance reference.
Distantly-supervised methods. First, we report distant supervision results as KB-Matching. We classify DS-NER methods into three groups. 1) DS-NER without Self-training consists of AutoNER Shang et al. (2018), Conf-MPU Zhou et al. (2022), RoBERTa-ES Liang et al. (2020). CuPUL is directly comparable with these methods. 2) DS-NER with Self-training includes BOND Liang et al. (2020), RoSTER Meng et al. (2021), SCDL Zhang et al. (2021b) and ATSEN Qu et al. (2023). These methods apply teach-student or training augmentation steps to further boost the DS-NER performance. 3)Span-based DS-NER models includes Top-Neg Xu et al. (2023), CLIM Li et al. (2023) and SANTA Si et al. (2023), and all other methods use sequence-based models. Previous work Li et al. (2023) shows that span-based NER models often outperform sequence-based NER methods in terms of effectiveness, albeit at the cost of increased algorithmic complexity. More details of baselines can be found in Appendix C.
5.2 Experiment Settings
We use the pre-trained RoBERTa as the backbone model for both the Voter and NER classifier [^3] We will release code upon paper acceptance.. For open-domain datasets, we use roberta-base [^4] https://huggingface.co/roberta-base. For bio-domain datasets, we use biomed-reberta-base [^5] https://huggingface.co/allenai/biomed_roberta_base. We report single-run results for the model performance and the random seed is set to 42. We employ PyTorch [^6] https://pytorch.org/ and conduct all experiments on a server with a Tesla A100 GPU (32G).
5.3 QTL Experiments
We discuss the performance of CuPUL on QTL dataset in this section.
5.3.1 Re-Examine Baseline Methods
In our exploration of baseline methods, we encountered several practical problems. For example, BOND [^7] https://github.com/cliang1453/BOND uses the test set for hyperparameter tuning during training, so we made modifications to their published code to tune hyperparameters on the validation set. RoSTER lacks tuning guidelines in the paper, so we used the reported hyperparameters from Meng et al. (2021) directly and only altered the max_seq_length to 300, since the sentence length of QTL is much longer. Finally, CLIM, Top-Neg, and SANTA lack runnable published code, so we did not include them in the evaluation. We also encountered some other practical issues, which are discussed in the Appendix E.
5.3.2 Results on QTL
Method QTL-strict QTL-relax
DS-NER without Self-training
KB-Matching 37.15 (82.95/23.93) 41.86 (93.46/26.97)
AutoNER 41.67 (69.07/29.83) 55.49 (83.17/41.64)
RoBERTa-ES 38.07 (76.30/25.37) 46.58 (91.15/31.28)
CuPUL-curr 54.75 (75.40/42.99) 62.94 (86.76/49.38)
CuPUL 58.02 (65.73/51.93) 70.24 (79.75/62.76)
DS-NER with Self-training
BOND 53.08 (60.89/47.04) 65.57 (77.97/56.57)
RoSTER 47.80 (73.12/35.51) 55.43 (91.35/39.79)
SCDL 43.62 (79.57/30.05) 50.18 (89.85/34.81)
ATSEN 46.23 (66.98/35.30) 51.64 (86.21/36.86)
CuPUL+ST 61.83 (58.65/65.38) 75.82 (72.96/78.91)
Table 1: Performance on QTL dataset: F1 Score (Precision/Recall) (in %). The best results are in bold, second best results are in underline.
Table 1 presents strict and relaxed precision, recall, and F1 scores for all methods on the QTL dataset. Note that CuPUL without curriculum learning (CuPUL-curr) is essentially equivalent to Conf-MPU when there is one entity type. KB matching reveals that QTL annotations suffer from low recall but have relatively high precision, creating a notable imbalance between these metrics. We observe that DS-NER baselines without self-training have limited recall improvement, resulting in weak performance. DS-NER baselines with self-training improve recall, but CuPUL and CuPUL+ST can further boost the recall, significantly outperforming all baseline methods. Specifically, strict F1 and relaxed F1 of CuPUL+ST outperform the runner-up by 8.75% and 10.25%, respectively.
5.4 Benchmark Experiments
We discuss the performance of CuPUL on existing benchmark datasets in this section.
5.4.1 Datasets and Metrics
Datasets: We conduct experiments on six existing benchmark datasets including CoNLL03 Liang et al. (2020), Twitter Liang et al. (2020), OntoNotes5.0 Liang et al. (2020), Wikigold Liang et al. (2020), Webpage Liang et al. (2020), and BC5CDR Shang et al. (2018). The first five are open-domain datasets, and BC5CDR is the bio-medical domain. More details and the statistics of these datasets are summarized in Appendix B.
Metrics: We use span-level Precision (P), Recall (R), and F1 scores as the evaluation metrics for all the datasets. These metrics require exact matches between predicted and actual entities. A continuous span with the same label is considered a single entity during inference.
Settings: For the benchmark dataset, we use small subsets of the validation set to tune the hyperparameters including learning rate, epochs, etc, to simulate the real-life DS-NER application scenarios. Detailed settings and statistics of the validation set can be found in Appendix D.
5.4.2 Results on Benchmark Datasets
Method CoNLL03 Twitter OntoNotes5.0 Wikigold Webpage BC5CDR
Fully Supervised
${}^{\#}$ 90.11 (89.14/91.10) 52.19 (51.76/52.63) 86.20 (84.59/87.88) 86.43 (85.33/87.66) 72.39 (66.29/79.73) 90.99 (${}^{\dagger}$
Span-based DS-NER models
${}^{\Diamond}$ 86.59 (86.25/86.95) - 69.72 (69.24/70.21) - 71.79 (78.40/66.72) 79.23 (81.74/76.88)
${}^{\Diamond}$ 80.55 (81.07/80.23) 52.86 (52.30/53.55) - - - 80.39 (82.09/78.90)
${}^{\Diamond}$ 85.4 (-/-) 53.8 (-/-) 69.6 (-/-) 70 (-/-) 67.9 (-/-) -
DS-NER without Self-training
${}^{\#}$ 71.40 (81.13/63.75) 35.83 (40.34/32.22) 59.51 (63.86/55.71) 47.76 (47.90/47.63) 52.45 (62.59/45.14) 64.32 (${}^{\dagger}$
${}^{\#}$ 67.00 (75.21/60.40) 26.10 (43.26/18.69) 67.18 (64.63/69.95) 47.54 (43.54/52.35) 51.39 (48.82/54.23) 79.99 (${}^{\dagger}$
${}^{\#}$ 75.61 (83.76/68.90) 46.61 (53.11/41.52) 68.11 (66.71/69.56) 51.55 (49.17/54.50) 59.11 (60.14/58.11) 73.66 (${}^{\dagger}$
${}^{\dagger}$ 79.16 (78.58/79.75) - - - - 77.22 (${}^{\dagger}$
CuPUL-curr 83.18 (83.69/82.68) 50.12 (47.48/53.07) 67.76 (65.66/70.00) 66.43 (58.89/76.18) 65.15 (62.89/67.57) 79.91 (75.07/85.43 )
CuPUL 85.09 (84.64/85.53) 54.34 (54.47/54.20) 68.06 (66.31/69.91) 70.53 (67.06/74.39) 73.10 (74.65/71.62) 81.57 (77.02/86.70)
DS-NER with Self-training
${}^{\#}$ 81.15 (82.00/80.92) 48.01 (53.16/43.76) 68.35 (67.14/69.61) 60.07 (53.44/68.58) 65.74 (67.37/64.19) -
[…]
${}^{\ddagger}$ 83.69 (87.96/79.82) 51.10 (59.87/44.57) 68.61 (67.49/69.77) 64.13 (62.25/66.12) 68.47 (68.71/68.24) -
${}^{\ddagger}$ 85.59 (86.14/85.05) 52.46 (62.32/45.30) 68.95 (66.97/71.05) - 70.55 (71.08/70.55) -
CuPUL+ST 86.64 (86.02/87.27) 54.78 (57.32/52.46) 68.20 (66.57/69.11) 70.19 (66.96/73.74) 74.48 (76.06/72.97) 80.92 (75.45/87.26)
Table 2: Performance on benchmark datasets: F1 Score (Precision/Recall) (in %). # marks the row of results reported by Liang et al. (2020). marks the row of results reported by Meng et al. (2021), where results for Twitter, OntoNote5.0 and Webpage are not reported in Meng et al. (2021). marks the row of results reported by Zhang et al. (2021b). $\Diamond$ marks the row of results from the method proposed paper respectively. $\dagger$ marks the results from Zhou et al. (2022). The best results are in bold, second best results are in underline.
Table 2 presents the overall span-level precision, recall, and F1 scores for all methods on benchmark datasets. Note that RoSTER was tested on a different version of the OntoNotes5.0 dataset Meng et al. (2021). Therefore, we exclude its reported results in Table 2 for a fair comparison. Note that the results of baseline methods are obtained from previous papers, and we did not rerun the experiments. The baseline methods used different validation sets for parameter tuning, which are much larger. We have the following observations.
Span-based DS-NER models. This group of baselines use span-based models. It is unfair to compare them with sequence-based models directly, so we just put it in Table 2 to inspire potential insights.
DS-NER Without Self-training. From Table 2, it is obvious that KB-Matching generally exhibits low recall and on four of the benchmark datasets, low precision as well. In contrast, noise-aware DS-NER models, like CuPUL, significantly outperform KB-Matching. This is confirmed in the table where CuPUL achieves the best F1 scores on four out of six datasets compared to all DS-NER models and displays competitive results on the OntoNotes5.0 dataset. On the OntoNotes5.0 dataset, almost all noise-aware DS-NER models have similar performances, implying that distant annotations may contain biases difficult for the models to address.
DS-NER With Self-training. The results for CuPUL+ST shown in Table 2 further verify that adding a self-training phase tends to enhance overall performance. When compared with DS-NER models that incorporate self-training, CuPUL+ST demonstrates superior performance on five datasets. This indicates that the CuPUL model benefits from the self-training approach, making it a versatile and effective tool for various datasets. Surprisingly, CuPUL and CuPUL+ST achieve higher F1 scores than the supervised baseline (RoBERTa) on the Twitter and Webpage datasets. This could be due to inconsistent distributions of training and test data for these datasets, and the supervised learning may have overfitted the training set due to its limited size.
5.5 Further Analysis
To further validate the effectiveness of CuPUL, we conducted additional analyses, including an ablation study. We are unable to use the QTL dataset for this purpose because the train set for QTL lacks ground truth, preventing us from analyzing the curriculum quality during training process. Specifically, we utilized open-domain datasets such as Wikigold and Twitter as representatives.
5.5.1 Difficulty Score Estimation
Figure 2: Distant Label Quality Token Level Positive Error Rate and Mean Difficulty Scores for Each Curriculum on Wikigold Dataset.
Figure 3: Distribution of the Difficulty Scores for Labeled Positives on Wikigold Dataset
For CuPUL, one assumption adopted is that difficulty scores can reflect the quality of distant supervision, where “easier” tokens have “cleaner” labels. To validate this assumption and evaluate the quality of the difficulty score estimation, we examine the correlation between the difficulty scores and the quality of distant labels. We use Wikigold as the testbed, and the results are illustrated in Figure 2.
For each training curriculum, we compute the token-level positive error rate (positive errors include false positives and positive type errors), and plot the rate for each curriculum using the left y-axis in Figure 2. We also compute the average difficulty scores for tokens in each curriculum shown with the right y-axis in Figure 2.
We can see that as the number of curricula increases, the average token difficulty scores and positive error rate have a clear increase. This illustrates a strong negative correlation between the difficulty scores and the quality of distant labels. Specifically, as the difficulty score increases, the quality of the distant labels significantly decreases. This result validates our assumption that “easy” data have cleaner labels and “hard” data have noisier labels. The clean data can initialize the model from a better starting point and improve the model’s robustness to noise in the latter curricula.
Another important assumption we adopt for the design of curricula is that the difficulty scores follow a long-tail distribution. We illustrate the distribution of difficulty scores estimated on the Wikigold dataset in Figure 3. It clearly demonstrates the long-tail phenomenon, with most tokens having low difficulty scores. This phenomenon can be observed in other datasets. Due to the space limit, we omit the plots for other datasets.
5.5.2 Ablation Study
Curriculum Learning To evaluate the effectiveness of curriculum learning in CuPUL, we compare it with two variations of itself. First, we use the five voters trained using positive and sampled negative examples and take the average of their soft label predictions as the result. The results are shown as voter ensemble in Table 3. Second, we include the result of CuPUL-curr from Table 2 since it is another variation. To evaluate the effectiveness of the Conf-MPU loss estimation for curriculum learning in CuPUL, we use the regular loss estimation, which considers unlabeled tokens as non-entity tokens, denoted as w/o Conf-MPU in Table 3.
Our analysis reveals the critical role of each component, as removing any of them results in a significant drop in the F1 score. Compared CuPUL-curr with w/o Conf-MPU, we find that CuPUL-curr consistently achieves higher recall. This is attributed to Conf-MPU primarily addressing false positives and partial false positives Zhou et al. (2022), leading to more tokens being predicted as entities, thereby enhancing recall. Conversely, w/o Conf-MPU exhibits higher precision since it tackles both false positives and positive type errors. Addressing positive type errors benefits both precision and recall, but the increase in precision is more pronounced compared to CuPUL-curr.
Method Wikigold Twitter
Precision Recall F1 Precision Recall F1
CuPUL 67.06 74.39 70.53 54.47 54.20 54.34
w/o Curriculum Learning
voter ensemble 56.88 74.88 64.65 35.52 49.52 41.37
CuPUL-curr 58.89 76.18 66.43 47.48 53.07 50.12
w/o Conf-MPU 59.31 75.86 66.57 58.91 47.04 52.53
Table 3: Ablation study on Wikigold and Twitter datasets. CuPUL is compared with variations without Curriculum Learning (voter ensemble only and Conf-MPU only) and without Conf-MPU loss in Curriculum Learning.
Distant Labels. In previous methods, a moderately well-trained model is often used to detect label noise, and the confidently predicted soft labels from the moderately well-trained model are often used to replace the noisy distant labels. Based on our previous experiments, the ensembled voters can be viewed as a moderately well-trained model, and the earlier curricula are formed with data that the moderately well-trained model can confidently predict. We study which labels should be used for curriculum learning in CuPUL, the voters’ ensembled soft labels or the noisy distant labels. Note that the ensembled labels used here are the soft labels of the voters’ ensemble. We use KL-divergence as the loss function in curriculum learning to learn from soft labels.
Figure 4 plots the results regarding F1 scores on test data with respect to incremental curriculum stages. We can see that CuPUL learns in almost all stages of the curricula, and the F1 value is steadily improving until the second last curriculum. However, using ensembled soft labels, the model has a good start but reaches the upper bound quickly. We have the following insights from this experiment. 1) A model that only learns from the confidently predicted labels and ignores the potential noisy data may converge faster but can be impacted by the performance bottleneck of the initial model. 2) the last curricula may contain high label noise, so training on the last curricula may degrade the performance slightly. However, thanks to the curriculum learning schedule, the model is overall robust to noise in the last curricula.
Figure 4: F1 scores of CuPUL on test data of Wikigold trained with Distant Labels (red) and Ensembled Labels from voters (blue) after each curriculum training stage.
6 Conclusion and Future Work
In this paper, we introduce a real-life DS-NER application in the animal science domain, named QTL, and we reveal the limitations of current DS-NER methods in practical DS-NER settings on QTL. To solve this issue, we propose a simple yet effective token-level curriculum-based PU learning (CuPUL) method, which strategically sequences the training data from “easy” to “hard”. Our experiments show that CuPUL not only mitigates the adverse effects of noisy labels but also achieves state-of-the-art DS-NER on many datasets. Through CuPUL, we demonstrate the effectiveness of curriculum learning in improving the performance of DS-NER systems.
Limitations
The "Baby Step" strategy in curriculum learning involves multiple repetitions of the first curriculum. Coupled with our power-law selector and curriculum scheduler, which tends to choose a larger initial curriculum, this may negatively impact efficiency if many curricula are established since the larger curriculum is repeatedly trained.
Ethics Statement
We comply with the ACL Code of Ethics.
Appendix
.1 Curriculum Learning
Curriculum learning was first proposed by Bengio et al. (2009) under the assumption that learning with reordering from “easy” samples to “hard” samples would boost performance. It has been applied in various applications, including neural machine translation Zhou et al. (2020); Platanios et al. (2019); Zhou et al. (2020); Wang et al. (2018), relation extraction Huang and Du (2019), reading comprehension Tay et al. (2019), natural language understanding Xu et al. (2020) and named entity recognition Jafarpour et al. (2021); Lobov et al. (2022); Wenjing et al. (2021).
Several studies aim to adopt curriculum learning philosophy for textual data and propose various difficulty-scoring functions and curriculum schedulers. Some methods measure sample difficulty with features derived from lexical statistics, e.g., sentence length and word rarity Platanios et al. (2019); Jafarpour et al. (2021), where longer sentences and rarer words are considered “hard”. Others use features from pre-trained language models Zhou et al. (2020); Wang et al. (2018); Liu et al. (2020). Most schedulers select samples with difficulty scores lower than a threshold Platanios et al. (2019). While Zhou et al. (2020) design a sample selecting function based on model uncertainty. Our approach, unique in applying token-level curriculum learning to DS-NER tasks, diverges from common sentence-level methods by utilizing Transformer-based models like BERT for context-aware token-specific predictions and gradient learning.
Appendix A Discussion of Loss Function
Two loss functions are popularly used for the DS-NER tasks. The first loss function is cross entropy (CE) loss:
[…]
where $f_{i,y_{i}}(x;\boldsymbol{\theta})$ is the prediction of token $x_{i}$ on class $j$.
Another commonly used loss function is mean absolute error (MAE):
$\mathcal{\ell}_{MAE}=|\boldsymbol{y}_{i}-f_{i,y_{i}}(x;\boldsymbol{\theta})|,$ (6)
where ||$\cdot$|| ⋅ | is L-1 norm of the vector and $\boldsymbol{y_{i}}$ denotes the one hot vector of $y_{i}$.
Comparing the two loss functions, $\mathcal{\ell}_{CE}$ is unbounded, and it grants better model convergence when trained with clean data (i.e., $y$ are ground truth labels) because more emphasis is put on difficult tokens. However, when the labels are noisy, training with the cross-entropy loss can cause overfitting to the wrongly labeled tokens. $\mathcal{\ell}_{MAE}$ is more noise-robust than $\mathcal{\ell}_{CE}$. It is bounded and treats every token more equally for gradient update, allowing the learning process to be dominated by the correct majority in distant labels. However, using $\mathcal{\ell}_{MAE}$ for training deep neural models generally worsens the convergence efficiency and effectiveness due to the inability to adjust for challenging training samples.
Considering the different characteristics of these two loss functions, in practice, we suggest using $\mathcal{\ell}_{CE}$ loss for tasks with more entity types and using $\mathcal{\ell}_{MAE}$ loss for tasks with fewer number of entity types.
Figure 5: Screenshot for online annotation tool teamtat.
Appendix B Datasets
To annotate the QTL dataset, domain experts use an online tool named teamtat [^8] https://www.teamtat.org/. The screenshot of the tool shows in Figure 5.
Here, we give a short description of the six benchmark datasets as follows:
  • CoNLL03 Tjong Kim Sang and De Meulder (2003) is built from 1393 English news articles and consists of four entity types: person, location, organization, and miscellaneous.
  • Twitter Godin et al. (2015) is from the WNUT 2016 NER shared task and consists of 10 entity types.
  • OntoNotes5.0 Weischedel et al. (2013) is built from documents of multiple domains like broadcast conversations, web data, etc. It consists of 18 entity types.
  • Wikigold Balasuriya et al. (2009) is built from a set of Wikipedia articles (40k tokens). They are randomly selected from a 2008 English dump and manually annotated with four entity types same as CoNLL03.
  • Webpage Ratinov and Roth (2009) comprises personal, academic, and computer science conference web pages. It consists of 20 web pages that cover 783 entities with four entity types same as CoNLL03 too.
  • BC5CDR comes from the biomedical domain. It consists of 1,500 articles, containing 15,935 Chemical and 12,852 Disease mentions.
The statistics of the baseline datasets are shown in Table 4.
Appendix C Baselines
Here, we give a short description of all the baseline methods: KB-Matching distantly labels the test sets using distant supervision, serving as a reference to illustrate the performance improvements given by other advanced DS-NER methods.
AutoNER Shang et al. (2018) trains the neural model with a “Tie or Break” tagging scheme for entity boundary detection and then predicts entity type for each candidate.
Conf-MPU Zhou et al. (2022) treats the NER task as a Positive-Unlabeled learning problem and utilizes the pre-learned confidence scores to enhance the model’s performance.
CLIM Li et al. (2023) addresses the imbalance problem in the high-performance and low-performance classes by improving the candidate selection and label generation.
SANTA Si et al. (2023) dealing with inaccurate and incomplete annotation noise in DS-NER by utilizing separate strategies.
Top-Neg Xu et al. (2023) selectively uses negative samples with high similarity to positives of the same entity type, improving performance by effectively distinguishing false negatives.
RoBERTa-ES Liang et al. (2020) trains a NER model using a RoBERTa-base model and adopts early stopping to prevent the model from overfitting to noisy distant labels.
BOND Liang et al. (2020) trains a RoBERTa model on distantly labeled data with early stopping and then uses a teacher-student framework to iteratively self-train the model.
RoSTER Meng et al. (2021) employs a noise-robust loss function and a self-training process with contextual augmentation to train a NER model.
SCDL Zhang et al. (2021b) conducts self-collaborative denoising with teacher-student framework. It trains two teacher-student networks, and the final reports come from the best model (teacher or student).
ATSEN Qu et al. (2023) develops a teacher-student framework with adaptive teacher learning and fine-grained student ensembling.
Appendix D Experiment Settings
For datasets CoNLL03, OntoNotes5.0, Webpage, Twitter, Wikigold, QTL and BC5CDR, the maximum sequence length is set as 150, 230, 120, 160, 120, 180, 280 respectively, to ensure the algorithm works correctly. Other parameters shows in Table 5. OntoNote5.0 and Twitter have more entity types, so we choose relatively small $\gamma$ values. We apply cross-entropy loss to OntoNotes5.0 and Twitter since they have more entity types and apply MAE loss to CoNLL03, Webpage, and Wikigold. For all the datasets, we train them with a batch size of 32 sentences and apply Adam optimizer Kingma and Ba (2014). For all the datasets, the number of voters $K$ and the number of curricula $C$ are set as 5 and 5, respectively. The curriculum selective factor $\tau$ is set to 0.5 for all the datasets. We use the same random seeds for all datasets.
To follow the real-life setting of DS-NER, we use the small subset of validation to adjust the learning rate on benchmark datasets. Other hyper-parameters are set according to data statistics (show in Table 4). For QTL dataset, baselines are reproduced using their published codes. Baselines CLIM, SANTA, and Top-Neg are excluded due to reproduction obstacles. Hyperparameters of baseline methods and CuPUL are tuned on the validation set.
Dataset Train Valid Test Types
CoNLL03 Sentence 14041 20 3453 4
Token 203621 475 46435
Twitter Sentence 2393 50 3844 10
Token 44076 719 58064
OntoNotes5.0 Sentence 115812 50 12217 18
Token 2200865 1090 230118
Wikigold Sentence 1142 20 274 4
Token 25819 579 6538
Webpage Sentence 385 20 135 4
Token 5293 120 1131
BC5CDR Sentence 4560 20 4797 2
Token 118170 533 124750
QTL Sentence 18706 21 1044 1
Token 514176 952 32251
Table 4: The statistics of involved DS-NER datasets, the valid set comprises a small subset from the original dataset, whereas the train set and test set utilize the entire original dataset.
hyper-parameter CoNLL03 Twitter OntoNotes5.0 Wikigold Webpage BC5CDR QTL
train set sentence # 14041 2393 115812 1142 385 4560 18706
voter drop negative 0.3 0.1 0.3 0.1 0.1 0.3 0.3
voter learning rate 1e-5 1e-5 1e-5 1e-5 1e-5 1e-5 1e-5
voter learning epochs 1 5 1 10 15 5 1
Conf-MPU $\gamma$ 20 10 20 10 10 20 20
curriculum learning stage epochs 1 2 1 2 2 1 1
curriculum learning learning rate 1e-5 7e-5 3e-5 1e-5 5e-5 1e-5 5e-5
Table 5: The hyper-parameters used in CuPUL
Appendix E Re-Examine Baseline Methods on QTL
We have explored various DS-NRE methods for QTL dataset. Our first attempt is AutoNER, which requires not only a dictionary for entity annotation but also a larger dictionary, called full-dict, for marking unknown labels, which leads to increased manual effort. To address this, we gathered a comprehensive dictionary of 26,620 potential trait entities. Unlike traditional machine learning approaches, AutoNER uses both a validation set and a test set during training and eliminates the need for hyperparameter tuning. In our exploration of RoBERTa-ES and BOND, we encountered the practice of using the test set for hyperparameter tuning during training. To rectify this, we modified the code to perform hyperparameter tuning on the validation set and conducted tests on the test set, focusing on hyperparameter tuning of early stop criteria and self-training period. For SCDL and ASTEN, we applied the hyperparameter tuning strategies outlined in the paper. Note that CuPUL without curriculum learning is essentially equivalent to Conf-MPU when there is one entity type. Therefore, Conf-MPU is not presented in the results.
Appendix F Parameter Study
Here, we perform parameter studies. Due to the simplicity of CuPUL, we mainly study two parameters: the number of voters $V$ and the number of curricula $\eta$. To ensure comparability of experimental results, we keep all other parameters fixed and only change the corresponding parameter ($V$ or $\eta$) to demonstrate their impact. The experiments are carried out on Wikigold.
Index 1 2 3 4 5 6 7 8 9
Token the regiment was attached to Howe ’s Brigade of $\cdots$
Ground Truth O O O O O ORG ORG ORG O
Distant Label O O O O O ORG ORG ORG ORG
Curriculum # 0 0 0 0 0 2 3 2 4
Index 10 11 12 13 14 15 16 17 18
Token the IV Corps of the Army of the Potomac
Ground Truth O ORG ORG O O ORG ORG ORG ORG
Distant Label O ORG ORG ORG O ORG ORG O O
Curriculum # 0 2 2 2 0 2 2 0 0
Table 6: Case study on Wikigold. The selected sentence is "After burying the dead on the field of Second Battle of Bull Run, the regiment was attached to Howe ’s Brigade of Couch ’s Division of the IV Corps of the Army of the Potomac where it replaced De Trobriand ’s 55th New York, Gardes Lafayette regiment on September 11, 1862." This table shows two pieces of this sentence.
F.1 Number of Voters $V$
Figure 6: Span Level Precision, Recall, and F1 scores of CuPUL with respect to Number of Voters $V$.
Figure 6 shows the effect of the number of voters $V$ to CuPUL performance. From the figure, we can see that when there are only two voters, the performance of CuPUL is poor. This is understandable because, with too few voters, the difficulty scores estimated are unreliable, which leads to a low-quality curriculum scheduler. As the number of voters increases, the performance of CuPUL also rapidly improves. When the number of voters is 4, it reaches a local maximum. Then, as the number of voters increases, the new voters can no longer provide new information for difficulty estimation, and the results of CuPUL are stabilized around 0.7. Therefore, with the consideration of computation efficiency, a moderate number greater than or equal to 4 can be chosen for the number of voters.
F.2 Number of Curricula $\eta$
Figure 7: Span Level Precision, Recall, and F1 scores of CuPUL with respect to Number of Curricula $\eta$.
Figure 7 shows the effect of the number of curricula to CuPUL performance. Like the number of voters, when the number of curricula is small, the performance of CuPUL is poor. Too few curricula can reduce the ability to distinguish between easy and difficult tokens, leading to ineffective curriculum learning. With the increase of $\eta$, the performance of CuPUL also improves and reaches the best performance at $\eta=5$. After that, as the number of curricula increases, the performance of CuPUL is relatively stable. The performance of CuPUL begins to decline after $\eta>8$. The decline may be caused by the data having been trained too many rounds, and the model starts to overfit to noisy labels.
Appendix G Efficiency Analysis
BOND RoSTER SCDL Conf-MPU CuPUL CuPUL-ST
Run Time 978s 2397s 4319s 732s 819s 1733s
16m18s 39m57s 71m59s 12m12s 13m39s 28m53s
Table 7: Efficiency analysis on CoNLL03, m means minute, s means second
In order to evaluate the efficiency of CuPUL, we undertook performance timing of the principal methods on CoNLL03, with the results displayed in Table 7. All tests were performed on an identical computing infrastructure. The training epochs for BOND and SCDL were preset to 5, while the parameter configurations for RoSTER adhered strictly to those detailed in their respective paper. The data in the table reveals that Conf-MPU had the least time requirement. Our approach, CuPUL, demonstrated competitive performance in this regard. Even when the self-training procedure was incorporated into CuPUL-ST, it maintained a substantial efficiency advantage relative to both RoSTER and SCDL.
Appendix H Case Study
To gain an intuitive understanding of how the curriculum helps CuPUL, we selected a sentence from the Wikigold corpus to show how CuPUL learns. As shown in Table 6, we give the tokens, ground truth labels, the distant labels, and the Number of Curricula for each token in the sentence. We assign each token an index for ease of discussion. We display a sentence in two lines and omit some repeated parts. As can be seen from Table 6, the two “of” (token 9 and token 16) are learned in different curricula. The one with the false positive label (token 9) is arranged in the fourth curriculum, whereas the one with the correct label (token 16) is learned early (the second curriculum). This shows that the pre-trained language model has the capability of providing prediction results for each token while retaining context information, and thus, the difficulty scores can be determined at the token level.